OpenStudy (anonymous):

what is the inverse of the function y=5x^3?

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

\[y=5x^3 \implies x = 5y^3\] solve for y now

OpenStudy (anonymous):

so would it y=5x/3?

OpenStudy (anonymous):

No where did you get the 3

OpenStudy (anonymous):

\[y^3 = \frac{ x }{ 5 }\] you will have to take the cube root not divide by 3.

OpenStudy (anonymous):

\[\huge x^{1/3} = \sqrt[3]{x}\] cube root

OpenStudy (anonymous):

\[\sqrt[3]{5x}\]

OpenStudy (anonymous):

so it would be that ^ correct?

OpenStudy (anonymous):

Nope, you take the cube root of both sides, \[\huge y^3 = \frac{ x }{ 5 } \implies y = \sqrt[3]{\frac{ x }{ 5 }} \implies f^{-1}(x) = \sqrt[3]{\frac{ x }{ 5 }}\]

OpenStudy (anonymous):

\[\large y^3 = \frac{ x }{ 5 } \implies y = \sqrt[3]{\frac{ x }{ 5 }} \implies f^{-1}(x) = \sqrt[3]{\frac{ x }{ 5 }}\]

OpenStudy (anonymous):

thank you !!